Mixed Bag: Some TCS testing, and the plan is for MODS and LBC checkout and filler science

High winds and some humidity - wind gusts are keeping the dome closed.

The TCS/SW group was here from the start and did some closed-dome testing.

Kellee launched a series of 50 50-sec exposures on LBC-Blue and LBC-Red to test code which should recover from RPC errors. But when they left, ~04:30 UT, there had not been any RPC errors. I'm continuing to run this while the dome is closed.

Started up LUCI and am running the flush_detector_once script. =NOTE: The remote workstation rm580h-1 (in front of polycom) has 4 screens, but the LUCI display works only on the middle horizontal one. On the two screens to the right of it (the one horizontal and one vertical) and on the left screen:
  • no images appear in skycat - its display is black,
  • nor in the LUCI display window - it is grey;
  • NXclient background is grey and icons are not visible

Started up MODS,
  • ran modsWake

LUCI: flush_detector hung on the 12th (of 150) exposures. RMGUI had hung. start_RMGUI.sh opened a new one, but it was hung also. From engineer account, stopped/restarted ReadoutManager service. Also restarted GEIRS. No joy. Finally noticed that the telescope service was hung. I restarted that. Now all seems good (followed Marcus' note in IT 3087).

MODS: sieveSnap.

05:50 AdSec problem on SX. 05:50:37.851 AOSL aos.AdsecSet.failed left AdsecSet[240] failed - IDL Dust detected in the mirror Called Doug who worked to get AdSec back up and running so it would set.

06:00 First sieveSnap image on red --- there is nothing in quadrant 1 (lower left)! See mods1r.20130923.0001.fits.
  • I repeated the sieveSnap and again see this mods1r.20130923.0002.fits. I believe there is still a red readout issue (IT #4813 will be updated). In 8K x 3K mode quadrant 1 is reading only bias.
  • 06:30 sieveSnap after quitting and restarting the IC on red. Still no response of quadrant 1 to light.

07:00 Collimating. Shell ripped. "COILS DISABLED" at 07:01:55.613.

07:30 Sent preset finally but error

 >> PRESET ACTIVE
** ERROR: PRESET IIF error - SetStars result status: Error, SetStars Error: target filter type invalid: , SetStars Error: target color type invalid:
** Abort, Retry, or Ignore? > a

** 5dice5_xy_0030+40_blue.img aborted with errors at line 12 command 'PRESET ACTIVE'
   ... clearing preset ...
DONE: clearstars Target and Guide Star coordinate data cleared...
Exit 2
This is a problem that we've run up against before (see email to nightreports from 18 May 2013 when we switched to 2013B; also new IT #4815). Now we have switched to 2013C but the earlier post-summer-shutdown MODS checkout was done on 20130905 with 2013B.

Called up Chris, who fixed it.

08:00 Sending a preset to test repeatability of offsets - 5 dice 5x on field at 00:30, +40. Seeing 1.4-2" on guider and we are at elevation 81 deg.
  • The seeing is up and down still. The seeing became so bad that an offset failed (no guide star found in guide box after resumeGuiding).
  • This may not be the best data to check offset repeatability, but using MODS has pointed out two issues: one resolved (IIF/MODS interface) and the other not yet (red channel quadrant 1).
  • On collimation - note that around the time of WFS image 112, the FWHM Mean was << FWHM guider, there was a CR in WFS image and the RMS WFE was ~ 1500 nm, even though we had been collimated earlier. What happened ?
  • When seeing is so bad, GCS server flashes yellow and reports "WARNING: guide star peak value is outside of desir"...The sentence is truncated.

09:00 Done with MODS 5dice. Switching to LUCI to observe some of the poor weather program of Rothberg. Seeing is still variable from 1.5- >2".
  • Shell ripped around the time Geno cancelled the MODS preset to switch to LUCI. Configured for LUCI, but shell is still being reset.

09:27 LUCI pointing and collimation. Seeing got as good as ~1.05", but bounced up to 1.8". The AZ-GRB project wants seeing better than 1", and Thuan has desired seeing ~1" also.

09:40 executeLUCIScript .sh HD_232766_M1Iab.acq (Rothberg). Seeing on guider ~1.5-2" and bouncing all over.
  • star saturated. Script used FeII + H.
    • switched to OH1060 + z (OK - #7).
    • Slit (z + clear - #8)
    • Run 4s x 15 DIT x NDIT script --> ~ 5000 counts (H=3.5)

  • executeLUCIScript.sh HD_20123_G5Iab.acq (OH1060+z and no offset).
  • executeLUCIScript.sh H-band_4x15_star_o2dcr.spec --> ~ 16000 counts per 4sec DIT (H=2.5)
  • NOTES on Scripts:
    • these are extremely bright stars Hmag =3.5-2.5 to <0
    • =acquisition scripts use FeII+H which is not sufficient. Need maximum blocking. Also offset, which is not needed for bright a star, well above sky background.
  • executeLUCIScript.sh HIP_15925_Telluric.acq
    • while acquiring, I realized that I had left the camera to N3.75 for shift image. I believe the previous 2 stars were still in the slits - the main offset was along the slit, and the X offset was small. Spectra were seen
  • seeing on guider during telluric is sub-arcsecond... looks like seeing is stabilizing.
  • run flush_detector script

10:45 Switch to LBCs.

11:14 Copoint/collimate at star 1300-0107625. RA05:00_DEC40:00
  • both red and blue require large -ve z4 (red: z4=-7631 and blue: z4=-3588).
  • red converged on 2nd iter with seeing 1.42" but blue still had large -ve z4.
* 11:23* lbcrangebal on 112343 and 112342
delta_IE = 0.9 arcsec and delta_CA = -8.31 arcsec
M1 SX : dX= -0.06mm, dY = 0.19mm, dRX = 4.08 arcsec, dRY = 1.32 arcsec
M1 DX : dX= -0.11mm, dY = 0.07mm, dRX = 1.59 arcsec, dRY = 2.45 arcsec

11:28 rotator center check
  • PA=0 (0.148 deg) r = 112836
  • PA=180 (-179.216 deg) r = 113223
  • PA=0 (0.826 deg) r=113434
  • PA=180 (-179.129) r=113633

11:37 dofpia, /x2
  • converged but with seeing ~ 1.4"
  • running superfoc7 at PA=0, 180 despite seeing, just in case there is no chance tomorrow.
  • Note PA change from 1st to 2nd image in sequence:
/newdata/lbcr.20130923.113905.fits[1]   0.94435 RB_rVfastextra
/newdata/lbcr.20130923.114044.fits[1]   0.18703 RB_rVfastextra

/newdata/lbcr.20130923.114322.fits[1]   0.21400 RB_rV_superfoc7
/newdata/lbcr.20130923.114404.fits[1]   0.37685 RB_rV_superfoc7
/newdata/lbcr.20130923.114444.fits[1]   0.36189 RB_rV_superfoc7
/newdata/lbcr.20130923.114529.fits[1]   0.38790 RB_rV_superfoc7
/newdata/lbcr.20130923.114609.fits[1]   0.34984 RB_rV_superfoc7
/newdata/lbcr.20130923.114649.fits[1]   0.38583 RB_rV_superfoc7
/newdata/lbcr.20130923.114730.fits[1]   0.39660 RB_rV_superfoc7

/newdata/lbcr.20130923.115248.fits[1]   -178.74435      RB_rV_superfoc7_pa
/newdata/lbcr.20130923.115329.fits[1]   -179.58704      RB_rV_superfoc7_pa
/newdata/lbcr.20130923.115411.fits[1]   -179.53008      RB_rV_superfoc7_pa
/newdata/lbcr.20130923.115451.fits[1]   -179.54547      RB_rV_superfoc7_pa
/newdata/lbcr.20130923.115531.fits[1]   -179.54353      RB_rV_superfoc7_pa
...
and blue does a similar thing:
/newdata/lbcb.20130923.113906.fits[1]   0.94177 RB_rVfastextra
/newdata/lbcb.20130923.114045.fits[1]   0.18117 RB_rVfastextra

/newdata/lbcb.20130923.114324.fits[1]   0.21472 RB_rV_superfoc7
/newdata/lbcb.20130923.114413.fits[1]   0.43071 RB_rV_superfoc7
/newdata/lbcb.20130923.114503.fits[1]   0.41671 RB_rV_superfoc7
/newdata/lbcb.20130923.114552.fits[1]   0.44262 RB_rV_superfoc7
/newdata/lbcb.20130923.114641.fits[1]   0.45382 RB_rV_superfoc7
/newdata/lbcb.20130923.114731.fits[1]   0.46907 RB_rV_superfoc7
/newdata/lbcb.20130923.114819.fits[1]   0.44704 RB_rV_superfoc7

/newdata/lbcb.20130923.115250.fits[1]   -178.73898      RB_rV_superfoc7_pa
/newdata/lbcb.20130923.115338.fits[1]   -179.49080      RB_rV_superfoc7_pa
/newdata/lbcb.20130923.115427.fits[1]   -179.46307      RB_rV_superfoc7_pa
/newdata/lbcb.20130923.115517.fits[1]   -179.45458      RB_rV_superfoc7_pa

12:00 Noted that only 4 RB_rV_superfoc7_pa180 appear on the blue side, but all 7 on the red side. Note a log warning for blue camera at this time:

12:00:30.08 CAMERA StatusPatch connection timeout [src/camera/camera.c:972]

Running FPIA now but no blue files are being written to /newdata. Blue does not report an error and blue, as well as red, has a timestamp when taking the exposure, e.g. lbcb.20130923.120915 and lbcr.20130923.120914. Blue appears to be working, but the data are not appearing in /newdata.

12:12 lbckill/lbcstart to see if that solves the problem with LBC blue data ... No, still no blue data.

In lbc log this problem seems to have started around 11:55 or 11:56 --- there is a message at 11:56:57: 11:56:57.778 N B CAMERA CCDCTRL issuing upload command aborted since file "lbcb.20130923.115605.fits" does not exist

12:19 Turning off/on LBCs. This did the trick... LBC is working again.

12:30 SkyFlats at Blank Field 1 RA 04:29:45.54 DEC +54 15 36
  • V/r --- a darkish splotch near the bottom left of chip 2, but nothing like what was seen on Friday
  • B/R --- same as above... reaching saturation so not many of these...
  • S/Y --- 124821/124842 OK but on blue 124900 the vignetting on the right seems to be a bit more than on the left
  • tried U/Y but saturated...

12:52 done. Geno closing.

13:00 At zenith, rotator 341 for flats for LUCI/Rothberg program. Changed scripts to set FLEXURE COMP = ON.
  • 13:09 filter wheel error. Filter wheel 2. Initialized and then set it manually to H, but no joy. Going to FW2 hardware failure instructions. Went through the hard reset. Clicked initialize on UI, and then tried again to manually set to H, but failed. Tried another hard reset, this time when I try to move the fw2 to clear, I get the error:
2013.09.23 13:21:35 s | instrument low | lucifer  | Lucifer ONE     | FilterUnit      | SequenceImpl.java#executeTransition(275)         | executing transition: "filter wheel two - move motor until state reached"
2013.09.23 13:21:35 s | instrument med | lucifer  | Lucifer ONE     | FilterUnit      | SequenceImpl.java#executeSequence(352)           | executing sequence: "filter unit - move negative sequence", parameter: {filter unit - filter unit: 102 position state: "represents the state of filter unit: 102 position: 4"}
2013.09.23 13:21:30 s | instrument med | lucifer  | Lucifer ONE     | FilterUnit      | SequenceImpl.java#executeSequence(424)           | finished sequence: "filter unit - initializing sequence"
2013.09.23 13:21:30 s | instrument low | lucifer  | Lucifer ONE     | FilterUnit      | SequenceImpl.java#executeTransition(318)         | finished transition: "filter wheel two - move motor until state reached"
Not going to continue for the flats/arcs... will have to resolve the fw2 problem first.

As for the darks: the ones for tonight are in 2013B_Fillers/Darks/rothberg_darks.cal (6s x 10 and 4s x 15, integrated, o2dcr) but will wait to do these until later since there was a lot of persistence.

Summary Not a very smooth night...
  • MODS1 red channel quadrant 1 not sensitive to light. (IT 4814)
  • MODS1 presets failed (IT 4815). We had this before - apparently presets are not compliant with TCS and a flag was set in 2013B to accomodate this. We called Chris and he set the flag in 2013C to continue this accomodation.
  • LUCI - observed 2 of Rothberg's stars --- neglected to change to N1.8 camera in "Shift Image" but the X offset onto the slit was small and the stars were there, albeit they could probably have been better centered.
  • LBCs -
    • obtained data for rotator center check.
    • ran superfoc7 sequences, albeit poor seeing, for PA=0 and PA=180. The final 3 blue images at PA=180 were not sent to /newdata. Had to stop/restart LBCs.
    • obtained a few flats at V/r, B/R and S/Y
  • LUCI - when sending cal script, FW2 failed. After two hard resets, I could not get it back to working and it is in this state.

-- OlgaKuhn - 23 Sep 2013
Topic revision: r3 - 23 Sep 2013, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback